OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 30 = x2 + x? (−5, 0) and (6, 0) (−5, 0) and (−6, 0) (5, 0) and (−6, 0) (5, 0) and (6, 0)

OpenStudy (anonymous):

at the points where the graph cuts the x-axis(x-intercepts), y=0 so, \[x ^{2}+x-30=0\]\[x ^{2}+6x-5x-30=0\]\[x(x+6)-5(x+6)=0\]\[(x-5)(x-6)=0\]\[x=5 or 6\] So, the last option is correct.

OpenStudy (anonymous):

(5,0)(0,6)?

OpenStudy (anonymous):

(5,0)(6,0)